OpenStudy (anonymous):

[6.06] Nick works two jobs to pay for college. He tutors for $15 per hour and also works as a bag boy for $8 per hour. Due to his class and study schedule, Nick is only able to work up to 20 hours per week, but must earn at least $150 per week. If t represents the number of hours Nick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario? t + b 20 15t + 8b = 150 t + b 20 15t + 8b 150 t + b 20 15t + 8b 150 None of the systems shown represent this scenario.

OpenStudy (e.mccormick):

Are these missing = signs for a reason?

OpenStudy (anonymous):

oh wait i wrote it wrong, one sec

OpenStudy (anonymous):

t + b ≥ 20 15t + 8b = 150 t + b ≤ 20 15t + 8b ≥150 t + b ≤20 15t + 8b ≤ 150 None of the systems shown represent this scenario.

OpenStudy (e.mccormick):

AH! That makes a lot more sense.

OpenStudy (e.mccormick):

"Due to his class and study schedule, Nick is only able to work up to 20 hours per week" What does that mean as far as the sum of the total hours go?

OpenStudy (anonymous):

thats how much he works a week

OpenStudy (e.mccormick):

So you are saying that "up to" is the same as "equals" ? Because that is the key here. What does "up to" mean in math. See my point there?

OpenStudy (e.mccormick):

In math we use: \(\ge\) Greater than or equal to \(\le\) Less than or equal to and = equal to (as well as others) But those words, "Greater than or equal to" and "Less than or equal to," there are other ways to say them. What would "up to" mean in that context?

OpenStudy (anonymous):

no idea

OpenStudy (whpalmer4):

I can spend up to 20 dollars means?

OpenStudy (whpalmer4):

Price <= $20 Price >=$20